three friends want to share a 20-kg bag of rice equally by weight. how many kilograms of rice should each friend get? write your answer as a mixed number


1
Expert's answer
2021-02-24T07:31:46-0500

"\frac{20}{3}=\frac{18+2}{3}=6\frac{2}{3}."